### v2

# Impact of the latest 22Ne+ reaction rates on nucleosynthesis in massive stars and galactic chemical evolution

Abstract
In massive stars (initial mass of 9 ), the weak (slow neutron capture) process produces elements between Fe and Zr, enriching the Galaxy with these elements through core-collapse supernova explosions. The weak s-process nucleosynthesis is driven by neutrons produced in the 22Ne()25Mg reaction during convective He-core and C-shell burning. The yields of heavy elements thus depend on the 22Ne()25Mg and the competitive 22Ne()26Mg reaction rates which are dominated by several narrow-resonance reactions. While the accuracy of these rates has been under debate for decades, recent experimental efforts, including ours, drastically reduced these uncertainties. In this work, we use a set of 280 massive star nucleosynthesis models calculated using different 22Ne()25Mg and 22Ne()26Mg rates, and a galactic chemical evolution (GCE) study to probe their impact on the weak -process elemental abundances in the Galaxy. The GCE was computed with the OMEGA+ code, using the new sets of stellar yields with different 22Ne+ rates. From GCE, we find that these rates are causing up to 0.45 dex of variations in the [Cu/Fe], [Ga/Fe], and [Ge/Fe] ratios predicted at solar metallicity. The greatest impact on the stellar nucleosynthesis and GCE results derives from uncertainties in the (,) strength () of the Ex=11.32 MeV resonance. We show that the variations observed in the GCE calculations for the weak -proess elements become negligibly smaller than dispersions found in observations once the is accurately determined within the uncertaintiy of 10–20% (typically reported experimental errors for the resonance) in future nuclear physics experiments.
1 Introduction
The slow neutron capture, namely the -process, produces about half of the solar abundances for elements heavier than Fe (Kappeler2011; lugaro:23). The other half of the heavy elements are mostly produced by the rapid neutron capture process (r-process Woosley1994; kajino:19; cowan:21; Siegel2022), with much smaller contributions from the p-process (rauscher:13; pignatari:16). The contribution of neutrino-winds ejecta in core-collapse supernovae (CCSN) (psaltis:22, and references therein) and from the i-process (e.g., Denissenkov2019; Choplin2019) is still a matter of debate.
The relevant stellar sources for the s-process elements are low- and intermediate-mass Asymptotic Giant Branch (AGB) stars and massive stars (initial mass 9 ). AGB stars produce the bulk of the solar s-process abundances between Sr and Bi (e.g., bisterzo:14; Prantzos2020), while massive stars are responsible for most of the production between Fe and Sr (weak -process; e.g., 2010ApJ...710.1557P). Most of the neutrons powering the s-process in AGB stars are generated by the 13CO reaction in the radiative 13C-pocket, with a smaller contribution by the 22NeMg reaction during Thermal Pulse events (e.g., Straniero2006; karakas:14; battino:16; lugaro:23). In massive stars, 22NeMg is the main neutron source powering the s-process, which is efficiently activated in the convective He-burning core and the convective C-burning shell (e.g., raiteri:91; baraffe:92; Kappeler1994; rauscher:02; the:07; 2010ApJ...710.1557P). The 13CO reaction may be activated in the C-burning core, but its products are typically further processed by more advanced stages and not ejected by the CCSN explosion (chieffi:98; Pignatari2013). Thus, the most relevant s-process stellar sources have been well-established. On the other hand, several uncertainties are affecting the stellar yields of s-process elements, mainly for CCSN. For these stars, quantitative predictions of theoretical stellar yields have been affected by uncertainties of both stellar models (e.g., Kappeler1994; 2016ApJS..225...24P; sukhbold:16; limongi:18; roberti:24) and nuclear reaction rates (2010ApJ...710.1557P; nishimura:17; Ota2021; Wiescher2023; pignatari:23, and references therein). The 22NeMg cross sections are particularly complicated because the reactions involve many resonances in their Gamow window. Therefore, their large uncertainties still have a significant impact on s-process nucleosynthesis (e.g., Wiescher2012; Adsley2021; Best2025). Additionally, as the 22NeMg reaction is an endothermic reaction, the neutron emission channels only open above a certain energy (=478 keV, Wang2021) where is a Center-of-Mass energy in 22Ne+ reaction). For resonances near the neutron threshold like this, -ray and neutron emission channels can compete. The n/ branching ratio of the resonance depends on the strength of its single-particle component (Wiescher2023). 22Ne in the massive stars can be consumed by 22NeMg reactions, leading to less active 22NeMg reactions in the end (Kappeler1994; Wiescher2012). As such, 22NeMg cross sections also need to be constrained to understand the s-process yields from the CCSN.
Like other charged-particle reactions, measuring these cross sections directly at stellar temperature is extremely challenging because of the small cross sections due to the extremely energy-dependent Coulomb barriers. The development of underground accelerator facilities, where neutron background environments are dramatically improved compared to those above the ground, is rapidly progressing (Best2016; Gao2022; Shahina2022). Nevertheless, the cross sections at low energy still need to rely on indirect measurements. Although both 22NeMg and 22NeMg cross sections have been studied by both direct (e.g., Wolke1989; Harms1991; Jaeger2001; Hunt2019; Shahina2024) and indirect (e.g., Giesen1993; Longland2009; deBoer2010; Massimi2012; Talwar2016; Massimi2017; Adsley2017; Adsley2018; Jayatissa2020; OTA2020135256; Ota2021; Chen2021) measurements for decades, many ambiguities remain.
Unambiguously determining these rates will drastically reduce uncertainties of nucleosynthesis calculations in massive stars, which are essential for improving the robustness of CCSN yields. The impacts of different 22Ne+ reaction rates on stellar yields are frequently discussed, but limited to specific stellar conditions (e.g., Longland2012; Talwar2016; Adsley2021; Ota2021). Their direct impacts on galactic chemical evolution (GCE) have not yet been studied. In particular, we expect that the s-process in massive stars is the main producer of solar abundances of Cu, Ga, and Ge; its contributions to other elements are less certain due to various uncertainties, including those in the processing of s-process rich layers by the CCSN explosion (2010ApJ...710.1557P; Prantzos2020).
In the present work, we therefore perform simulations to study the GCE of Cu, Ga, and Ge and other elements (from Zn to Sr), particularly focusing on their sensitivity to the 22Ne + reaction rates in massive stars. Different 22Ne + rates, including a conventionally referenced rate (Longland2012, hereafter LO12) and the rate we obtained from the experiment by OTA2020135256 (hereafter OT20), are used in our simulations. Since many uncertainties remain in the reaction rates, multiple reaction rates that adopt different 22Ne + cross sections have been published (e.g., Adsley2021; Wiescher2023). OT20 produces the smallest weak s-process stellar yields among these rates, (Ota2021) because the adopted rate for the () reaction is the smallest and that for the () reaction is relatively large. In contrast, the LO12 rate produces the largest weak -process yields among the four reaction rates mentioned here. Because the 22Ne + reaction in -process stars is dominated by resonance reactions, GCE sensitivity to 22Ne+ rates can be investigated by varying the strength of individual resonances. We estimate the impact of individual resonances on the GCE using this approach, comparing the GCE by LO12 with that by OT20. We aim to determine a desirable precision and upper limits for future measurements to constrain the strengths of the relevant 22Ne+ resonances. The same stellar set provided by Ritter2018 is used in our nucleosynthesis calculations for the GCE simulations. We thus focus on studying the impact of different reaction rates on the stellar abundances and GCE.
The paper is organized as follows. The present uncertainties in 22Ne reaction rates and the strengths of the relevant resonances are discussed in Section 2. We describe the stellar evolution models used in our study and summarize the present knowledge of the production of light -elements: Cu, Zn, Ga, and Ge in Section 3. Presented in Section 4 are the main features of the stellar yields and GCE models used in this work. In Sections 5, the calculated GCEs for Cu, Zn, Ga, and Ge from the pre-supernova (SN) explosions are compared with observations, discussing the sensitivity to the varied strengths of individual resonances and possible impacts by the SN explosions. Finally, summary and conclusions are discussed in Section 6.
2 22Ne+ reaction rates
Since the ground states of both 22Ne and have spin-parity () of 0+ and their reaction Q-value for populating the compound nucleus, 26Mg, is highly positive (Q = 10.614 MeV), 22Ne+ reaction primarily proceeds through resonant capture to natural-parity states in the highly excited 26Mg (see review by Wiescher2023). While the compound nucleus’s level density is high, only a limited number of resonances that have natural-parity and high -cluster strength are expected to contribute to the 22Ne+ reaction. The expectation has been verified by direct measurements, also confirming these resonances’ widths () are narrow and do not overlap between each resonance. For a given resonance, the key properties that determine its contribution to the stellar rate are the 22Ne+ resonance energy (Ex = + ECM) and the resonance strengths, and . Two (one possible) resonances, at =11.32 and 11.17 MeV, have been intensively studied due to their importance in the weak s-process’s temperature (0.3–0.35 GK). Although the weak s-process is reactivated at about 1 GK, the final abundances are less sensitive to the current uncertainties of the resonance strengths at the relevant energy.
One of the two important resonances is well-identified at ECM = 0.702 MeV (Ex = 11.32 MeV). Both the () and () stellar rates are dominated by this resonance. It has been observed in both direct () (Harms1991; Drotleff1991; Giesen1993; Jaeger2001; Shahina2024) and () experiments (Wolke1989; Hunt2019; Shahina2022). Its is well established with these published measurements in excellent agreement with each other (Best2025). In contrast, extracted from the direct () measurements are in poor agreement, making the for the Ex = 11.32 MeV resonance the outstanding contributor to uncertainties on the total 22Ne()25Mg stellar rate.
At lower temperatures (0.25 GK), both reaction channels may be dominated by one or more resonances between the neutron threshold (=11.09 MeV) and ECM 0.635 MeV (Ex 11.25 MeV). In this region, in particular the resonance at Ex 11.17 MeV has long been controversial (see e.g., Longland2012) since reported by Berman1969. Jaeger2001 searched for this state in their direct () measurement, setting an upper limit on the at Ex 11.17 MeV, 60 neV. Recently, Shahina2022 also observed no strong transitions around this energy region in their () experiment, setting an upper limit of 150 neV on a candidate resonance at Ex 11.17 MeV. Additionally, the for the natural-parity resonances identified by Massimi2012, the lowest-lying resonance above , = 11.11 MeV, remain poorly constrained (60 neV) by direct measurements (Jaeger2001). Although this resonance has never been reported to have strong -cluster strength by indirect measurements, it may have the potential to completely dominate the () rate below 0.3 GK due to its proximity to the Gamow peak (Ota2021). While the = 11.32 MeV and 11.17 MeV resonances are critical for 0.2–0.35 GK, two resonances below the neutron threshold (therefore no () cross sections), 11.08 and 10.95 MeV resonances, could contribute to consume 22Ne via () reactions at 0.1–0.2 GK (OTA2020135256) before () reaction is activated at a later (hotter) stage of the star (Kappeler1994; Wiescher2023).
These situations are well-described (in e.g., OTA2020135256; Adsley2021). In this work, we will therefore study the sensitivity of the GCE to the strengths of these individual resonances ( and ), including those above 11.32 MeV, by varying their strengths.
3 Stellar models and production of the elements copper, zinc, gallium and germanium
3.1 Stellar models and CCSN explosions
For the nucleosynthesis calculations presented in this work, we used the 20 massive star models from the stellar set by Ritter2018 111Initial mass =12, 15, 20, and 25 (where ZAMS means Zero Age Main Sequence) and metallicity = 0.0001, 0.001, 0.006, 0.01, and 0.02 calculated with the MESA code, rev. 3372 (2011ApJS..192....3P). The CCSN models are calculated with a semi-analytical approach based on the Sedov blast wave (SBW) solution for the determination of the peak velocity of the shock throughout the stellar structure (Sedov1946), while the remnant mass is determined using the delayed explosion prescription from Fryer2012. For further detail, readers are referred to 2016ApJS..225...24P, Ritter2018 and Roberti2024.
For =0.01 and 0.02, the initial abundances adopted are solar-scaled (Grevesse1993). At = 0.006 and below, -enhanced abundances were used, where the isotopes of 12C, 16O, 20Ne, 24Mg, 28Si, 32S, 36Ar, 40Ca, and 48Ti are enhanced compared to a solar scaled composition (see Table 1, in Ritter2018), based on stellar observations. Different -enhancements are applied to different species, varying between [Ca/Fe]=+0.222 (where the enhancement is applied to 40Ca, and the remaining Ca isotopes are solar scaled), and [O/Fe]=+0.886 (the enhancement is only applied to 16O). Although the solar abundances by Grevesse1993 have been updated several times (more recently, see Magg2022), these variations would not affect the results of the present study.
The Multi-zone Post-Processing Nucleosynthesis in Parallel code (MPPNP) developed by the NuGrid collaboration is used for the calculations of the complete nucleosynthesis (e.g., 2016ApJS..225...24P) as used in Ritter2018. The MPPNP post-processes the production and destruction of all relevant isotopes during stellar evolution and during CCSN explosions, allowing us to derive stellar isotopic yields. In this work, the nucleosynthesis simulations are performed for the aforementioned stellar models, where only the 22Ne+ rates are modified. The stellar structure and CCSN models can be used unchanged for different 22Ne+ rates, as the reactions do not impact energy generation and therefore do not affect stellar evolution properties. Both the pre-SN and post-SN nucleosynthesis were calculated using version V2.2 of the JINA Reaclib database if reaction rates from experiments were not available. For instance, for the 22Ne(,n)25Mg rate and the 22Ne(,)26Mg rate, Ritter2018 adopted Jaeger2001 and Angulo1999, respectively. Experimental neutron capture reaction rates were taken, when available, from the KADoNIS compilation (Dillmann2006). We refer to 2016ApJS..225...24P and Ritter2018 for more details of all the reaction rates used for the post-processing calculations.
3.2 Productions of Cu, Zn, Ga, and Ge
While the bulk of the s-process products are expected to be built in pre-SN hydrostatic evolution, the shock wave and explosive nucleosynthesis during CCSN explosions can affect the ejected s-process yields (e.g., rauscher:02). Therefore, explosive nucleosynthesis should be taken into account to derive the complete stellar yields. In the following part of this sub-section, we will explore the productions of the elements Cu, Zn, Ga, and Ge, which are located just above the iron-group element Ni, and are produced by the s-process in massive stars (e.g., raiteri:91; raiteri:91a; rauscher:02; the:07; 2010ApJ...710.1557P). For the remainder of this work, we will focus our analysis and discussions on these elements. As mentioned, it is established that most of the solar Cu, Ga, and Ge are produced by the s-process in massive stars, while the origin of Zn remains a puzzle (e.g., 2010ApJ...710.1557P; Prantzos2020; kobayashi:20). Thus, it is beneficial to discuss the production processes of these elements in this subsection to better understand their productions in the different layers of the CCSN ejectas, which will lead to better understanding of the impact of the 22Ne+ rates discussed in the later sections.
3.2.1 Copper
In the solar system, Cu is made up of 69.15% of 63Cu and 30.85% of 65Cu. In Figure 1, we show the abundance profiles for 63Cu in the CCSN ejecta of the 15M⊙ at =0.02 and =0.0001 models calculated by Ritter2018. Note that here denotes the abundance in mass fraction of a given isotope for each mass coordinate in the model222In a given differential mass zone , at the mass coordinate , .. Although the two models have the same initial mass, the model at lower shows a more relevant primary333i.e., not directly affected from the initial composition of the star, in contrast to secondary. explosive nucleosynthesis component, within an -rich region in the deepest 0.2-0.3 M⊙ region of the ejecta, right above the newly formed proto-neutron star (PNS), where the -process takes place (e.g., woosley:92; 2016ApJS..225...24P). Therefore, we will discuss the primary production of Cu (and other elements discussed in this subsection) in CCSN using the =0.0001 model.
In the model at =0.02, 63Cu shows a mild s-process production in the material ejected from the former convective He shell at mass coordinates between 3.5 M⊙ and 4.5 M⊙. 63Cu at 3.1 M⊙–3.5 M⊙ is produced by n-process (meyer:00; pignatari:18, and references therein) mostly as radiogenic444By radiogenic production, we mean that the given stable isotope, in this case 63Cu, is not ejected directly as itself but in the form of radioactive isotopes that will decay later to 63Cu, based on their decay half-life. 63Ni and 63Co, triggered by the 22Ne(,n)25Mg activation during explosive He burning. The bulk of the s-process production is visible at about 1.8 M⊙–3 M⊙ as ashes of the pre-explosive C burning. Traces of more extreme explosive components are present in the most internal ejected layers, but are only barely noticeable in this model.
In the model at =0.0001, the 63Cu production discussed above is about three orders of magnitude smaller, due to the smaller of the progenitor and the well-expected reduction of both the neutron source 22Ne (as it is secondary) and the iron seeds (e.g., baraffe:92; raiteri:92; pignatari:08; roberti:24). Instead, a strong primary production is obtained in the -rich freezout material at 1.6 M⊙–1.8 M⊙. 63Cu show abundances up to a few per mill in mass fractions, and is mostly produced from the radiogenic contribution from 63Zn, with much smaller direct production as 63Cu and the radiogenic contribution from 63Ga.
In Figure 2, the same production pattern as Figure 1 is observed for 65Cu, which are ejected in the same CCSN layers as 63Cu. The s-process makes 65Cu mostly directly, with a marginal radiogenic contribution from 65Zn (which is made from neutron-capture on 64Zn). The n-process during the explosion instead produces 65Cu mostly as 65Ni, with a small contribution from 65Co. 65Cu also has a strong primary component above the PNS in the model at low , specifically as radiogenic contribution from 65Zn and 65Ga (Figure 2, right panel).
3.2.2 Zinc
The solar Zn composition is made up of 49.17% of 64Zn, 27.73% of 66Zn, 4.04% of 67Zn, 18.45% of 68Zn and 0.61% of 70Zn. The nucleosynthesis of Zn in massive stars has been studied extensively for both CCSN and the most energetic hypernovae (kobayashi:20, and references therein), due to the puzzling high abundance of Zn with respect to Fe observed in the early Galaxy. This is mostly due to the primary production of 64Zn (and partly 66Zn). In our models, an effective s-process production is obtained for all Zn isotopes except for 64Zn. Since the s-process marginally produces Zn, here we focus in detail on 67Zn (Figure 3) and 68Zn (Figure 4). Analogous figures for 64Zn, 66Zn and 70Zn are available in Appendix.
In Figure 3 left panel (=0.02), 67Zn shows a less efficient s-process production compared to 66Zn and 68Zn. This is a well-known effect in the s-process, where the neutron capture rate of 67Zn (with an odd number of neutrons) is much larger than the nearby species with an even neutron number. Interestingly, we can see in this model a strong n-process radiogenic production of 67Zn mostly as 67Ni and 67Cu, which in the integrated stellar yields is comparable to the s-process contribution. In the model at =0.0001 (Figure 3 right panel), there is a strong -process radiogenic production as 67Ge and 67Ga. The nucleosythesis of 68Zn shown in Figure 4 is very similar to the Cu isotopes. At low , the primary 68Zn production is dominated from the radiogenic contribution from 68Ge.
3.2.3 Gallium
In the solar system, Ga is made of 69Ga (60.108%) and 71Ga (39.892%). The nucleosynthesis of Ga in CCSN shows remarkable similarities to Cu. Due to the difficulty in observing Ga in stars, a good understanding of their nucleosynthesis production could allow us to use Cu as an observational diagnostic for Ga (2010ApJ...710.1557P).
In Figure 5, 69Ga is produced directly by the s-process. In the left panel, the n-process component is built by the radiogenic contribution of 69Ni, 69Cu and 69Zn. In the right panel, the -process component is made of radiogenic 69Ge and small amounts of 69As.
3.2.4 Germanium
The stable Ge isotopes 70Ge, 72Ge, 73Ge, 74Ge and 76Ge account for 20.52%, 27.45%, 7.76%, 36.52% and 7.75% of the solar Ge, respectively. 70Ge is defined as an s-only isotope, meaning that its production is due to the s-process only, as the r-process contribution is shielded by the stable isobar 70Zn (e.g., Kappeler2011). In Figure 7, left panel, we confirm this also for the n-process, where there is no direct or radiogenic contribution feeding 70Ge abundances. On the other hand, such a definition may not be entirely accurate, since it is possible to have a primary 70Ge production by the -process, mostly directly making the isotope and with some small radiogenic contribution from 70Se and 70As. A small primary component for Ge is indeed observed in metal-poor stars (cowan:05; 2010ApJ...710.1557P).
Unlike 70Ge, in Figure 8, left panel, 74Ge can also be ejected with a small n-process component due to the radiogenic contribution of 74Zn. However, the direct s-process production will dominate the production of this isotope. In Figure 8, right panel, 74Ge does not show relevant primary production via the -process, due to the shielding from the stable isobar 74Se.
Analogous figures are available in the Appendix for 72Ge, 73Ge and 76Ge.
4 Stellar yields and GCE calculations
4.1 Nucleosynthesis and stellar yield table - testing the impact of the 22Ne+ rates
As mentioned, we performed the post-processing nucleosynthesis simulations for the 20 massive star progenitor models in the stellar set considered, at the given initial mass ( = 12, 15, 20, and 25 ) and ( = 0.0001, 0.001, 0.006, 0.01, and 0.02). For each progenitor, we performed three calculations using the upper bound, recommended, and lower bound for each of the LO12 and OT20 rates. The upper bound of each reaction rate was probed by using the upper limit of their 22Ne(,n)25Mg rate and the lower limit of their 22Ne(,)26Mg rate. The recommended calculation was made using the recommended rates for both reactions. For the lower bound, we instead used the lower limit of their 22Ne(,n)25Mg rate and the upper limit of their 22Ne(,)26Mg rate. A total of 120 (20 stellar models 6 reaction rates) complete nucleosynthesis simulations were thus performed for LO12 and OT20, including the CCSN nucleosynthesis, to compare with the original yields set by Ritter2018.
Additional pre-SN nucleosynthesis calculations are performed for six variants of the OT20 recommended rate, and for the more recently evaluated rates by Adsley2021 and Wiescher2023 to better interpret the differences observed between the LO12 and OT20 rates. Thus, an additional 160 nucleosynthesis simulations were performed.
Figure 9 shows the post-SN abundance profiles of 65Cu and 70Ge for the 15 at =0.02 model produced in the present work (similar to Figure 2 and 7 left panels). As expected from the discussion in subsection 3.2, large variations caused by the different 22Ne+ rates are observed mainly between 1.8 and 3 (where the -process is active), for both 65Cu and 70Ge. The largest variation is seen between the high bound setup of LO12 rates and the low bound setup of OT20. Such variations are mostly generated in the convective He-burning core in the progenitor star, carried in the following shell C-burning regions and finally ejected by the CCSN explosion. Regarding the explosive processes discussed earlier, the n-process is not largely affected by 22Ne+ uncertainties, since it operates at temperatures around 1 GK (e.g., pignatari:18) and variations between different rates are much smaller than those induced by under He-burning conditions (see e.g., 65Cu in Figure 9).
The process is also not affected by the 22Ne+ rates. On the other hand, uncertainties in the explosion model may significantly affect the appearance of this component, and the use of a more sophisticated CCSN explosion simulation is mandatory to study in detail the nucleosynthesis nearby the PNS (see, e.g., curtis:19; boccioli:23; BR24, and references therein). We plan to address this challenge in future work.
In the following section, we focus on the impact of the pre-SN yields on the GCE, aiming to isolate the effects of uncertainties in the 22Ne+ reaction rate. This approach allows us to disentangle these effects from those related to CCSN modeling, which will affect the production of Zn and the other elements discussed in the present work more at low .
4.2 GCE Simulations
We calculated stellar yields from all the 280 nucleosynthesis simulations described in Section 4.1 and used them as inputs for GCE calculations. Stellar wind yields, i.e., the time-integrated mass lost by stellar winds throughout the lifetime of the AGB and massive stars, are included in the total yields used for GCE calculations, although their contribution to the elements considered in the present work is negligible.
For AGB stars (with initial masses 1-7), we used the original yields from Ritter2018, while for Type Ia SN (SNIa), we used the fast deflagration W7 yields from Iwamoto1999. We did not perform additional simulations on these astrophysical sources because the productions of Cu, Zn, Ga, and Ge are dominated by massive stars. Likewise, we did not consider the contribution from popIII stars or Hypernovae. These sources could potentially contribute to the abundances of these elements at very low (e.g., kobayashi:20). For the present work, we will focus on the GCE of stars with [Fe/H]-2. Stellar yields are interpolated below and above the transition mass of 8 for the AGBs and massive stars, respectively (see Ritter2018syg, for detail).
We used a two-zone GCE code, OMEGA+ (2018APJ_Cote). The first (inner) zone consists of the star-forming region in the galaxy modeled by OMEGA (2017APJ_Cote) and the second (outer) zone represents the circumgalactic medium (CGM). The mass fraction of a given isotope () in the gas (with its mass denoted by ) within the inner zone at a given time will be obtained within OMEGA+ as follows (Cote2016; 2017APJ_Cote). The temporal evolution of the gas within the inner zone () is calculated with the following equation:
where represents the gas inflow from the CGM at after the Big Bang (=0), the stellar ejecta, the gas used for star formation, and is the gas outflow into the CGM. Similarly, the evolution of gas within the CGM () is calculated as:
where and are the gas coming in or out from external intergalactic medium (IGM), respectively. In the present work, and are set to 0. We tested different constant values for them following the prescription in 2017APJ_Cote (and references therein; ) and within the range listed in Table 1 of 2018APJ_Cote () and observed negligible impacts on our GCE results.
At each time-step , OMEGA creates a simple stellar population (SSP) using SYGMA (Ritter2018syg) to model the stellar ejecta. The SSP is a group of stars created at the same time with the same initial metallicity (, zero-age main sequence), where the initial mass function (IMF) of 10.1046/j.1365-8711.2001.04022.x defines the mass distribution of the AGB and massive stars. The stellar ejecta, , at a certain time is calculated by taking the sum of the SSP’s ejecta from =1 to the total number of SSPs that have formed by time . Thus,
where depends on the initial mass population, , the initial metallicity, (=), and the age of the population, .
For other variables (, , and ), we followed a prescription provided by 2019APJ_Cote, in which model parameters were tuned to best reproduce observational constraints for the Milky Way disk (kubryk:15).
We used a two-infall model of gas inflow by 1997ApJ...477..765C, which accounts for two episodes of exponential gas inflow. This model is described with the following equation:
where the time of maximum infall is set by the , the continuation time of the infall is set by and for the first and second episode, respectively, and the normalization constants and represent the strength of each infall.
| Parameter | Value |
| 46 | |
| 5.9 | |
| (Gyr) | 1.0 |
| (Gyr) | 0.8 |
| (Gyr) | 7.0 |
| 0.23 | |
| (Gyr) | 1.0 |
| 0.52 |
The stellar formation rate (SFR) of the inner zone, , is set with the following equation:
where is the stellar formation efficiency and is the star formation time scale. The outflow of the inner galaxy, , is proportional to the SFR, where the scale factor is set using a mass loading factor, :
The values of the parameters used above are taken from 2019APJ_Cote and 10.1093/mnras/stac3180, and they are listed in Table 1.
In Figure 10, we plotted some important galactic properties: (SFR), (the gas mass of the galaxy), , , the CCSN rate, and SNIa rate as a function of the galactic age. All these diagnostics fall within the observational limits, reproducing the results from 2018APJ_Cote. Note that the two infall-episodes are not as clearly distinguishable in these plots (e.g., ) as those of 1997ApJ...477..765C. The bimodality derived from the two episodes is, however, more clearly visible when using different parameter values as shown in 2019APJ_Cote. Furthermore, since we do not include the threshold in stellar formation process as 1997ApJ...477..765C adopted, the two episodes are more smoothly connected to each other, making the bimodality less apparent.
It should also be noted that, as discussed in Cote2016 and 2019APJ_Cote, choice of different and has non-negligible impact on GCE. For instance, when using sets of the parameter values to reproduce the high and low limits of the observational constraints (see Fig. 10 and 2019APJ_Cote), variations of dex for Cu and Zn and dex for Ga and Ge are found in the GCE results below [Fe/H]–0.5, respectively. However, we confirmed that the relative differences observed in GCE using different reaction rates, which are the main interest of the present work, are negligibly influenced by the choice of the parameter sets. No GCE variations are indeed observed near solar .
Since the values of , the mass-loading constant that controls , are not largely varied in the above parameter sets (0.52, 0.50, and 0.45 for the best-fit, low, and high limits, respectively; defined in Table 1 of 2019APJ_Cote), we further tested different values for to specifically see the role of in the GCE. The resulting differences in the GCE curves are negligibly small except for the final (or [Fe/H]) where the GCE reaches at the end of the simulation (=13 Gyrs). About 0.1 dex higher (lower) [Fe/H] is reached with a lower (higher) value of =0 (1.0) because the enriched heavy elements in the Galaxy are exchanged for less enriched medium in the CGM. Note that we stop the simulations at =13 Gyrs. The same simulation time was assumed e.g., in Kobayashi2006, and using the more recently reported Milky Way galaxy’s age (13.6 Gyrs) only leads to small differences in our GCE results.
As additional benchmarking, we plotted the temporal evolution of [Fe/H] and in Figure 11. In the figure, reaches the solar value (=0.014 from Asplund2009) by the time the solar system was formed (8.5 Gyrs) and continues toward moderate supersolar enrichment (=0.02). [Fe/H] however is slightly underestimated (–0.2 dex) at 8.5 Gyrs, possibly because we used -enhanced initial abundances for stellar yield calculations at 0.006. Since the initial abundances of Fe are relatively small for the corresponding (see above), it could lead to a smaller amount of Fe ejecta from massive stars and SNIa (see e.g., Ritter2018syg). We thus conclude that our GCE simulations are reasonably calibrated and uncertainties derived from the calibration process do not impact the conclusions of this work.
Further details of the GCE code and the adopted GCE setup can be found in 2018APJ_Cote.
4.3 Observational data
The simulated GCE was compared with observational data collected from STELLAB: a code with a database of spectra observations from individual stars within the Milky Way (see 2017APJ_Cote). Most of STELLAB’s data are from the galactic disc, halo, and globular clusters within the Milky Way. In this work, we use the available stellar observations from the Milky Way disk to compare with our GCE simulations.
To supplement the STELLAB data for Ga and Ge, we imported stellar abundance data from JINAbase (2018ApJS..238...36A) and added observations from Ivans2003, and from cowan:05, Sneden1998, and Placco2015 for Ga and Ge, respectively.
5 GCE Results
5.1 Impact of the 22Ne+ rate uncertainties on the GCE of Cu, Zn, Ga and Ge: simulations using preSN yields
In Figure 12, we show the pre-SN GCE curves for Cu, Zn, Ga, and Ge calculated from the massive star yields using different 22Ne+ reaction rates, in comparison with observational data. Abundance evolution for each element is shown as [X/Fe] ratios555We remind that [X/Y] = (mass fractions are used for and ). The solar values are from Asplund2009. Only the gas component (i.e., isotopes not trapped within stars) is considered.
The GCE curves shown in Fig. 12 are produced using the yields from the nucleosynthesis calculated with LO12 and OT20 rates including their uncertainties. Additional curves are calculated with other 22Ne+ rates, in particular the recommended rates from Adsley2021 (Adsley-TAMU, hereafter AD21), from Wiescher2023 (Notre-Dame (ND), hereafter ND23), and the original Ritter2018 configuration (hereafter CR18). Although not shown in the figure, we confirmed that the upper and lower limits of the GCE curves using the yields from ND23 and AD21 rates lie within the curves bounded by the LO12 and OT20 rates.
| Element | Cu | Zn | Ga | Ge | As | Se | Br | Kr | Rb | Sr |
| 29 | 30 | 31 | 32 | 33 | 34 | 35 | 36 | 37 | 38 | |
| LO12-OT20 | 0.18 (6) | 0.22 (7) | 0.41 (12) | 0.44 (14) | 0.53 (17) | 0.37 (15) | 0.38 (14) | 0.37 (19) | 0.41 (20) | 0.41 (26) |
A wealth of observational data is available for Cu and Zn; therefore, these elements have been studied in detail in previous GCE research (e.g., Matteucci1993; Timmes1995; 1997ApJ...477..765C; Mishenia2002; Bisterzo2005; Kobayashi2006; mishenina:17; kobayashi:20; lach:20). The availability of observational data for Ga and Ge is instead scarce, with only a handful of metal-poor stars measured in the range considered for this work ([Fe/H]).
The [Cu/Fe] and [Zn/Fe] variations due to the OT20 and LO12 uncertainties are less than 0.05 dex, while they increase up to 0.1 dex for [Ga/Fe] and [Ge/Fe] (see Figure 12). However, the OT20 and LO12 distributions do not overlap in the figure, with [Cu/Fe] and [Zn/Fe] variations between the two rates (LO12 upper bound and OT20 lower bound) up to 0.4 dex at solar metallicity, and up to about 0.7 dex for [Ge/Fe] and [Ga/Fe].
The [X/Fe] values relative to the GCE calculation performed with yields using other rates typically evolve between the OT20 and LO12 limits. Except for Zn, all simulations are consistent with the solar values within the observational uncertainties. The inconsistency found in Zn is not surprising because of the strong primary Zn component that is not accounted for in these simulations (2010ApJ...710.1557P). As mentioned earlier, the source of the primary Zn production is still a matter of debate (Bisterzo2005), with possible large contributions from hypernovae in the early Galaxy (kobayashi:20, and references therein). Thus, the Zn discrepancy shown in Figure 12 is not related to the current sensitivity study of the 22Ne reaction rates.
While the solar abundance of Cu is better reproduced by the LO12 or the CR18 rates, those of Ga and Ge seem to be better reproduced by the CR18 rate, lying between the OT20 and LO12 rates. Future stellar observations of Ge and Ga for typical of the Milky Way disk would be needed to derive stronger constraints. For Cu, the GCE models tend to underproduce [Cu/Fe] for lower than solar. This could be justified by the missing primary Cu component in the pre-SN yields.
The variation of [X/Fe] obtained using the OT20 and LO12 recommended rates is summarized in Table 2. Here we report the variation for all elements between Cu and Sr. In general, 0.2-0.5 dex variation was found for [X/Fe]. However, the impact on the GCE of As and other heavier elements should be re-evaluated in the future since the massive star’s -process contribution for As and other heavier elements is expected to be smaller compared to AGB stars and/or explosive nucleosynthesis (Kappeler2011).
5.2 Resonance Sensitivity
In this section, we investigate the difference in the LO12 and OT20 rates in more detail. We study the impact on the GCE of the differences in the two rates, focusing on the contributions from individual resonances. The of relevant resonances to the weak -process identified as having large uncertainties by Ota2021 and by the recent review article Wiescher2023 are summarized in Table 3. It should be noted that the adopted strengths and upper limits (denoted by ””) have substantially different meanings in stellar reaction rates as described below. The upper limits are not calculated using the specified .
While the adopted strength was used in the stellar reaction calculations, the upper limit was used to determine the median and probability distribution assuming the Porter-Thomas distribution (Longland2010; Longland2012) in the present work. As such, the actual (e.g., median) strength used in stellar rate calculations is much smaller than the upper limit used as it is in e.g., Angulo1999; Kappeler1994; Massimi2017. For example, for the =11.17 MeV resonance in LO12 has a median of 10-8 eV ( 4.3010-7 eV upper limit). Thus, the resonance’s impact on the stellar yield is negligibly small using either the LO12 or OT20 (5.4210-11 eV) rate.
| (MeV) | 10.95 | 11.09 | 11.11 n | 11.17 | 11.32 n | 11.32 n |
| (MeV) | 0.336 | 0.471 | 0.498 | 0.551 | 0.702 | 0.702 |
| (LO12) | 8.7010-15 | 0 | 5.8010-8 | 4.3010-7 | 1.4010-4 (30) | – |
| (OT20) | 9.0010-14 (270) | 2.8510-10 (157) | 1.1010-9 | 5.4210-11 | 4.2110-5 (69) | – |
| (Case 1) | 1.4010-4 | |||||
| (Case 2) | 6.6010-7 | |||||
| (Case 3) | 1.3510-8 | |||||
| (Case 4) | 0 | |||||
| (Case 5) | 8.7010-15 | |||||
| (Case 6) | 0.5 |
As mentioned in Section 4, we performed GCE calculations for six different variants of the OT20 recommended rate (Case 1-6, as summarized in Table 3).
In Case 1, the LO12 rate was used for the =11.32 MeV resonance.
In Case 2, the for the =11.17 MeV resonance was increased to 660 neV, which is obtained by Talwar2016 from their indirect measurement. Note that the existence of the resonance was invalidated by the latest direct measurement experiment by Shahina2022, and an upper limit of 150 neV was determined. Thus, the median strength should be much smaller (10 neV). Still, we use 660 neV to probe the maximum impact of this resonance.
In Case 3, we increased the for the =11.11 MeV resonance to 13.5 neV (approximately the value adopted by Massimi2017; see Ota2021). It should be noted that the resonance is dominated by (,n) channel and the is constrained to the same level as the =11.17 MeV resonance (60 neV) by direct measurements (Jaeger2001). The upper limit used by LO12 (58 neV) reflects this value. We studied this resonance only for its potential contributions, as theorized by Massimi2017. However, since this is the lowest possible resonance for () channel, the weak -process could be activated at lower temperature (0.3 GK) if the resonance strength is high.
In Case 4, the for the =11.09 MeV resonance was set to 0 as LO12.
In Case 5, the same for LO12 was used for the =10.95 MeV resonance (but not as an upper limit).
In Case 6, we decreased the OT20’s (,) reaction rate at 1 GK by a factor of 2. Since a number of resonances (primarily 11.8 MeV and above) contribute to the 22Ne(,)25Mg reaction, the relevance of the 22Ne(,)26Mg becomes negligible. In this temperature range, we estimated the summed impact of all the high energy resonances by simply applying a correction factor to the reaction rate.
Results of the GCE calculations are shown in Figure 13, which are normalized with respect to the OT20 rate GCE calculation. Apart from the coordinates around [Fe/H]-1, the differences from the OT20 curve are nearly independent of the metallicity. The change in GCE observed by each case is summarized in Table 4. For all elements considered, the differences in the GCE by the LO12 and OT20 rates are due mostly ( 90%) to the differences of the at =11.32 MeV resonance (Case 1).
Given the difference in at =11.32 MeV is only a factor of 3.3, it is clear that this is the most impactful on s-process nucleosynthesis in massive stars among the remaining uncertain in the 22Ne + reaction. The recent direct measurement by Shahina2024 reported the to be 100 (22) eV, which varies from OT20 by a factor of 2.4. Its impact on the GCE is therefore estimated to be 0.15 dex (Cu and Zn) and 0.3 dex (Ga and Ge), which would result in elemental abundances distributed between OT20 and LO12. Upcoming verification of their results by future measurements in underground facilities can considerably reduce the uncertainties in the GCE (Masha2022; Best2025). If is unambiguously constrained with a precision of better than 10–20% (or e.g., =10-20 eV; which are typical or slightly better experimental uncertainties reported for this resonance), it will reduce the uncertainties from this resonance on the GCE simulations to 0.05 dex for Cu and Zn, and 0.1 dex for Ga and Ge, respectively.
As mentioned above, while the =11.17 MeV (Case 2) and 11.11 MeV (Case 3) resonances vary the GCE as much as the =11.32 MeV resonance, the used in this exercise does not correspond to the used in LO12. However, the GCE variations observed in Case 2 and 3 indicate the upper limits that future nuclear physics experiments should probe, without assuming any probability distributions (e.g., Porter-Thomas). Since such assumptions are ambiguous depending on the probability distribution, it is more instructive to determine the upper limit such that the resonance is negligible if its upper limit is below a certain value, especially when the limits may be reachable by the experiment.
For the of the =11.17 MeV resonance (Case 2), the GCE varies by 0.2–0.4 dex when increasing the from 0 to 660 neV. Assuming a simple linearity, we could safely say that the GCE uncertainties will be 0.05 dex for all the elements considered if the upper limit was reduced 10 times from 660 neV. If future experiments set an upper limit of 60 neV on this resonance (the current limit: 150 neV by Shahina2022), rather small GCE uncertainties (0.05 dex) would be achievable. If the for the =11.11 MeV resonance (Case 3) was similarly constrained by reducing the upper limit (13.5 neV) 10 times, the GCE uncertainties would be 0.05 dex for all the elements considered. Therefore, a 1 neV upper limit would be enough for this resonance. Again, no experimental evidence has been reported that the =11.11 MeV resonance has a large -cluster strength. In addition, the direct measurements show no signature of a strong resonance near the corresponding energy. Nevertheless, it will be beneficial to constrain the upper limit of this to 1 neV in future direct measurements due to its potential impact and experimental feasibility. At underground facilities such measurements can be possible (Rapagnani2022). It should also be mentioned that these upper limits may slightly change as the neutron productions compete among different resonances using limited amounts of 22Ne. For instance, if the dominant =11.32 MeV resonance has a higher value, the impact of the lower energy resonances could be reduced. Thus, even higher upper limits on the resonances may be acceptable.
Case 4 and 5 do not affect the GCE results, indicating for these resonances have already been well constrained. Case 6 has a negligible impact on the GCE of the elements shown in Figure 13. However, the production of neutron-rich isotopes close to branching points in the C-burning shell could be affected, due to neutron density variations.
| Case | 1 | 2 | 3 | 4 | 5 | 6 | Default |
| [Cu/Fe] | +0.15(5) | –0.15(5) | +0.14(5) | +0.01(5) | +0.01(5) | 0.00(5) | 0.18(6) |
| [Zn/Fe] | +0.20(4) | –0.09(4) | +0.18(4) | +0.01(4) | +0.01(4) | +0.01(4) | 0.22(7) |
| [Ga/Fe] | +0.38(9) | –0.18(9) | +0.33(9) | +0.03(9) | +0.03(9) | 0.00(9) | 0.41(12) |
| [Ge/Fe] | +0.40(8) | –0.16(4) | +0.36(4) | +0.03(4) | +0.03(4) | 0.00(4) | 0.44(14) |
5.3 Post-SN GCE
The purpose of the present work is to study the impact of uncertainties of the 22Ne+ rates on the s-process in massive stars; we have explored this topic in great detail in the previous sections. As we have seen in Section 4.1, the -process component found in some CCSN models by Ritter2018 is not affected by the 22Ne+ rates, and depends instead on the CCSN explosion setup. It is possible to roughly estimate the primary CCSN production of Cu to 6-16%, based on observations of metal-poor stars (2010ApJ...710.1557P). A similar contribution could be estimated for Ge (around 20%) and Ga, but the limited number of spectroscopic observations for these elements limits the ability to derive relevant constraints.
Using the same GCE framework adopted for the previous simulations, we have calculated a GCE model using the CCSN yield table provided by Ritter2018. Figure 14 shows the GCE curves of Cu, Zn, Ga, and Ge. As a reference, the pre-SN GCE scatter between the LO12 upper limit and OT20 lower limit (see section 5.1) is shown as green bands, with the pre-SN GCE using the CR18 rate. Within the range covered in the figure, the GCE model using CCSN yields for Ga and Ge does not appear to diverge significantly from the GCE models adopting the pre-SN yields, with the few observations available consistent with the results.
Simulations of Cu and Zn show instead some tension with observations (e.g., mishenina:17). The [Cu/Fe] trend is consistent with the solar abundances, but it is quite flat for different . Observations show instead solar ratios between ⊙ and [Fe/H] -0.5, but then they start to decrease for lower . Using the same phenomenological considerations in 2010ApJ...710.1557P, this would imply that in the GCE model adopting Ritter2018 CCSN yields, about 30-50% of the solar Cu would be made by primary CCSN components, assuming that 1/3 or 1/2 of the solar Fe is provided by massive stars. Therefore, in Ritter2018 standard CCSN yields the primary Cu production is overestimated by a factor between 2 and 8 with respect to observations, and this is directly noticeable in the CCSN abundance profiles from 12 and 15 stars at 0.006 (see e.g., Figures 1 and 2 for the 15 M⊙ model at Z=0.0001). An overestimation compared to observations is also obtained for the [Zn/Fe] ratios. This is interesting, since Zn has been historically under-produced by GCE simulations, leading to scenarios where a boosted Zn contribution from hypernovae has been claimed at low (kobayashi:20, and references therein). Our GCE simulations have no hypernovae, but we still see the huge potential impact of the -process on Zn, from the most internal 0.1-0.2 M⊙ in the CCSN ejecta.
More careful investigation is in progress, using observations of these elements (Cu, Zn, Ga, and Ge) in metal-poor stars, to find new constrains for the deepest CCSN ejecta. However, this will not affect the s-process production of these elements.
6 Summary and future work
While 22Ne(,)25Mg and 22Ne(,)26Mg are among the most important reactions for s-process nucleosynthesis, their reaction rates still have significant uncertainties. Jayatissa2020; OTA2020135256 and Ota2021 used indirect measurements to report weaker strengths of relevant resonances, such as the =11.32 MeV for the (,n) channel and 11.17 MeV for the (,) channel. The most recent direct measurement by Shahina2024 obtained a more moderate value of for the =11.32 MeV resonance which is more consistent with other experiments within their experimental uncertainties (22%). Their measurements of the for the =11.17 MeV are consistent with our measurements, implying that this resonance probably does not contribute significantly to -process nucleosynthesis. As described in Wiescher2023, there are remaining uncertainties in for other resonances below and above the energies of these two resonances.
The uncertainties from each of these resonances affect abundance yields of theoretical stellar models and the GCE of the s-process elements. This work provides the first detailed and quantitative estimation of the impact of different 22Ne+ rates on the s-process in massive stars and on the GCE of its main nucleosynthesis products. In particular, we focus on the production of the element group Cu, Zn, Ga, and Ge with Z = 29–32. For stellar nucleosynthesis simulations, we used the post-processing nucleosynthesis code, MPPNP, developed by the NuGrid Collaboration. For GCE simulations, we used the GCE code, OMEGA+. For the analysis, we considered 14 different 22Ne+ reaction rate combinations. For each reaction rate configuration, the nucleosynthesis in 20 massive star models (four initial stellar masses, 12, 15, 20, and 25 M⊙, and five metalicities, Z=0.02, 0.01, 0.006, 0.001, and 0.0001) were calculated, for a total of 280 nucleosynthesis models.
In comparison with observations, the simulated GCE reproduce their solar abundances except for Zn. Our simulations underproduce Cu at low partly because of the uncertainties of the GCE parameters used in the present work. The stellar models that are used can also impact the obtained final yields. For instance, pumo:10 discusses that convective overshooting during the He core burning influences how much 22Ne is available and the duration of neutron-capture conditions. Their work demonstrates that using different values for the overshooting parameter (10-5–0.035; c.f., 0.022 used (in the present models Ritter2018)), the weak abundances could decrease/increase by a factor of 2–3 in 15–25 M⊙ stars. Furthermore, fast rotating massive stars may potentially increase the process yields compared to the non-rotating stars used in the present work, especially at low (pignatari:08; frischknecht:16; roberti:24). As well as () cross sections, uncertainties of the other nuclear cross sections relevant to neutron economy in the process stellar environment such as neutron poison (e.g., 16O()17O) could also help explain the inconsistency with the observations. The ratio of 17O()/17O() cross sections influences the available neutrons for the weak process. As shown in best:13, the ratio’s current uncertainties could vary the abundances of weak elements by an order of magnitude. Therefore, many uncertainties in the physics input remain that hinder comparisons between the GCE simulations and observational data. Nonetheless, we consider the relative differences in GCE by different 22Ne reaction rates reported in the present work are still valid.
We obtain that using our recent 22Ne(,)25Mg and 22Ne(,)26Mg rates, the elements Cu, Zn, Ga, and Ge result in up to a factor of 1.5-3 (0.2-0.45 dex) reductions in their galactic mass fractions compared to GCE models using the rates by Longland2012. Such reductions are almost exclusively explained by the decrease of at the =11.32 MeV resonance. It confirms that this resonance is undoubtedly the current dominant source of the 22Ne+ rate’s uncertainties for the s-process in massive stars.
We also investigated the impact on GCE predictions to relevant lower (10.95, 11.08, and 11.11 MeV) and higher energy resonances (11.32 MeV), which could contribute to the s-process in massive stars during convective He-core burning and during convective C-shell burning, respectively. Apart from the =11.11 MeV resonance, these do not impact the GCE when reasonable or upper limits are assumed.
As such, it is critical to unambiguously constrain the of the =11.32 MeV resonance with a precision better than 10–20% in future experiments. Thanks to the worldwide ongoing/planned experimental efforts at underground nuclear physics facilities such as LUNA, this should be feasible within the next few years. Lowering the upper limits on the of the =11.17 MeV and of the =11.11 MeV to negligible levels (60 neV and 1 neV, respectively) will also be possible (Rapagnani2022). These experiments will have 10-100 times lower background environment compared to the work by Jaeger2001. We have demonstrated that well-constrained 22Ne reaction rates will result in much smaller uncertainties than the observational data. This will provide a crucial milestone to better constrain stellar evolution models of massive stars and the chemical enrichment history of our Galaxy.
Finally, we discussed the GCE of Cu, Zn, Ga, and Ge using CCSN yields. The primary explosive components observed in metal-poor stars are found in our nucleosynthesis simulations. Such nucleosynthesis products are located in the deepest 0.1-0.2 M⊙ ejected, and they depend on the CCSN explosion setup. We show that they are not affected by the 22Ne reaction rates, or by their current uncertainties.
We emphasize the importance of including complete CCSN yields to compare stellar and GCE simulations with observational results more consistently, as they complement the pre-SN nucleosynthesis production. In future work, we indeed expect to use the s-process elements discussed here to constrain CCSN models by comparisons of GCE and stellar results with galactic archaeology observations, applying similar calibration methods adopted in the past using Ni (limongi:18) or Zn (Nomoto2013) as reference elements. In turn, we will apply the newly-constrained CCSN models to study GCE for all the elements.
prin_2022rjlwhn. We acknowledges support from the ChETEC-INFRA project funded by the European Union’s Horizon 2020 Research and Innovation programme (Grant Agreement No 101008324). We thanks access to viper, the University of Hull HPC Facility. This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group (http://csa.phys.uvic.ca) at the University of Victoria and hosted on the Computed Canada Arbutus Cloud at the University of Victoria. SO was supported in part by the IReNA visiting fellowship. This work benefited from
interactions and workshops co-organized by The Center for Nuclear astrophysics Across Messengers (CeNAM) which is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0023128.
We thank Prof. Michael Wiescher and Dr. Richard deBoer (Univ. Notre Dame) for providing us with their reaction rates. We also thank Dr. K. A. Womack (Univ. Hull) for providing us with OMEGA+ parameters.